You wish to split the numbers [1,2,3,4,5,6,7,8] into two sets of 4 numbers, such that the following 2 conditions are satisfied.
• The sums of the numbers in both sets are equal
• The sums of the squares of the numbers in both sets are equal
I have found the two sets of nymbers [1,4,6,7] and [2,3,5,8] but i cannot find a relationship between the sets that will enable me to split the original list.
Help please !!

Answer by dabanfield(803)

You've actually solved the problem. The numbers in the two sets need to add up to a common total and their squares also need to add to a common total:
Set A = {1,4,6,7}
Set B = {2,3,5,8}
Sum of A's elements is 1+4+6+7 = 18
Sum of B's elements is 2+3+5+8 = 18
Squares of A = {1,16,36,49}
Squares of B = {4,9,25,64}
Sum of A squares is 1+16+36+49 = 102
Sum of B squares is 4+9+25+64 = 102